OpenStudy (anonymous):

Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form. A. (y + 1) = 3(x – 2) B. (y – 1) = 3(x + 2) C. (x + 2) = 3(y – 1) D. (x – 2) = 3(y + 1)

OpenStudy (anonymous):

basically, all you need is \[(y-(y intercept))=m(x-(x intercept))\] where m is the slope.

OpenStudy (anonymous):

Ive been working on this for and hour now i still cant get it! I would really appreiate it if you would just tell me.

OpenStudy (anonymous):

y intercept is y=1 x intercept is x=-2 m is slope m =3 plug that in to the equation I just gave you and VOILA.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

hwo do i plug t in?

OpenStudy (anonymous):

y=mx+b 1=3(-2)+b Solve for b, which will give you your y-intercept. :o)

OpenStudy (anonymous):

@Butterfly16 look at the answer format. @Cookie_009 there is no t

OpenStudy (anonymous):

Can you just plzzzzz tell me or show me how to plug it in?

OpenStudy (anonymous):

REPLACE yintercept with 1 and xintercept with -2 and m with 3!! I don't think that's too difficult.

OpenStudy (anonymous):

slope-point form: Point Slope form: y-y1=m(x-x1) Point Slope Form: y-1=3(x+2) Does that make sense?

OpenStudy (anonymous):

Im gonna kill myself!!!!!!!!!!!!

OpenStudy (anonymous):

@Butterfly16 just gave the answer -_-

OpenStudy (anonymous):

Is it a b or c or d

OpenStudy (anonymous):

y-1=3(x+2) Match it yourself.